OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 B. -12

OpenStudy (samgrace):

\[gradient = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] we know all of these except \[x_1\] \[(x_1,y_1) = (x, 6)\] \[(x_2, y_2) = (10, -3)\] gradient is equal to -3/2 so.. \[-\frac{ 3 }{ 2 } = \frac{ -3 - 6 }{ 10 - x }\] rearranging: \[x = 6\]

OpenStudy (samgrace):

Sorry, rearranging gives x = 4

OpenStudy (anonymous):

thanks.